# PathoHR: Hierarchical Reasoning for Vision-Language Models in Pathology

### v1

@@ 1 Introduction
Recently, vision-language (VL) models have gained substantial advances (Goel et al., 2022; Li et al., 2022), fostering an integration of computer vision and natural language processing across a wide range of domains (Wei et al., 2024; Das et al., 2024). In medical image analysis, contrastive learning-based approaches have been extensively employed to align large-scale medical images with corresponding diagnostic reports, thereby enabling zero-shot classification and grading without additional fine-tuning (Xie et al., 2024; Phan et al., 2024). This paradigm markedly reduces the reliance on high-quality annotated data and enhances the scalability of medical imaging models. Nevertheless, in contrast to other imaging modalities such as X-ray and MRI, pathological image analysis presents distinct and more formidable challenges due to its higher structural complexity and subtle visual cues of the images.
Figure 1: Radar charts for compared models on PathoHR-Bench.
Pathological biopsy images remain the gold standard for tumor diagnosis, as they capture intricate cellular structures and morphological details (Ruiz López et al., 2017). However, their subtle intra-class variations pose significant challenges for disease classification. Existing approaches predominantly rely on task-specific models tailored to individual applications, such as Gleason grading of prostate cancer and breast cancer subtyping (Pati et al., 2023; Bulten et al., 2022). In contrast, pathological texts are inherently fine-grained, characterized by specialized medical terminology, precise pathological reasoning, and structured diagnostic logic. These narratives not only describe subtle visual cues but also offer critical insights for diagnostic grading and prognosis prediction.
@@ 2 Motivations
The lack of hierarchical reasoning capabilities significantly limits the reliability and clinical interpretability of existing VL models in pathological analysis. A further limitation lies in current evaluation practices, which predominantly rely on standard retrieval or classification metrics and offer limited insight into models’ structural understanding or reasoning ability. While recent studies in the natural image domain have proposed benchmarks for evaluating compositional reasoning in VL models (Zhao et al., 2022b; Diwan et al., 2022), these approaches are not directly applicable to pathology due to fundamental domain differences in textual structure and semantics. Descriptions in natural image datasets are typically open-domain and emphasize visually salient features. In contrast, pathological texts are "closed-domain", relying on highly specialized medical knowledge with much tighter semantic relationships and long-range logical reasoning chains.
@@ 3 Pathology Hierarchical Reasoning Benchmark (PathoHR-Bench)
At the text perturbation level, we have designed three core tasks to simulate the challenges encountered in real-world pathology texts, including information loss, semantic shift, and order variation:
i). Assessing sensitivity to information loss: We employ saliency-driven phrase deletion by leveraging the Pathology Language Image Pretraining (PLIP) (Huang et al., 2023) model to compute the similarity between each phrase and the corresponding image. The two most salient pathological terms are removed to create adversarial text variants, simulating the model’s reliance on critical textual cues and assessing its behavior under different deletion orders. It reflects how the model allocates information between visual and textual inputs while evaluates its robustness to local information losses.
ii). Assessing sensitivity to semantic drift: We employ LLM-guided token substitution, where a randomly selected word in each text is masked and then predicted using a pretrained medical BERT model (BioBERT) (Kenton and Toutanova, 2019; Lee et al., 2020). The generated adversarial samples test the model’s robustness to semantic drifts and fine-grained semantic comprehension. It evaluates whether the VL model can accurately differentiate between subtly different but diagnostically significant textual descriptions.
@@ 4.1 Pathology-guided textual perturbation
A simple but effective approach of negative text generation is to apple a collection of pre-defined linguistic rules to match and replace words associated with specific entity types or patterns (Doveh et al., 2023). We pre-defined seven pathological attribute dimensions (pathological states and grading, morphological features, histochemical characteristics, staining methods, anatomical structures and organs, biomolecular features, color information) via PubMed (Namata et al., 2012) and MeSH (Lipscomb, 2000), followed by random substitution at the level of pathological attributes to generate controlled perturbations.
@@ 4.5 Loss Function
The total positive loss $L_{pos}$ is computed as:
$L_{pos}=L_{pos}^{text_{T}}+L_{pos}^{text_{I}}+L_{pos}^{text}$ (9)
Finally, the full finetuning loss of our proposed method can be written as:
$L=L_{con}+\alpha\cdot L_{neg}+\beta\cdot L_{pos}$ (10)
5 Experiments
5.1 Datasets
The ARCH dataset (Gamper and Rajpoot, 2021) is the only widely available pathological image-text paired dataset, comprising 8617 pairs extracted from pathology textbooks and PubMed research articles. We compared the performance of our proposed method with current pathological VL methods, including baseline CLIP (Radford et al., 2021), PLIP (Huang et al., 2023), BiomedCLIP (Zhang et al., 2023), CONCH(Lu et al., 2024), QuiltNet(Ikezogwo et al., 2023) and MI-Zero (Lu et al., 2023c) on PathoHR-Bench. For a fair comparison, all methods were trained on ARCH dataset (except for baseline CLIP) and tested under the same testing splits and inference prompts. For the zero-shot task, we utilized six publicly available datasets covering a range of cancer types, including four datasets at the patch level and two at the whole-slide level. Detailed descriptions of these datasets are provided in the Appendix B, and implementation details can be found in the Appendix C.
5.2 Comparison of existing VL models
In contrast, our proposed method effectively leverages limited pathology-specific data to enable fine-grained compositional reasoning while maintaining strong performance on conventional classification tasks. As shown in Tables 1 and 2, the model achieves consistent improvements across nine textual perturbation settings and six zero-shot classification benchmarks. Notably, substantial gains are observed on CRC100K and UHU (with fine-grained cancer subtypes), while more moderate gains are seen on PanNuke and DigestPath (with coarser tumor/normal classification). These results further support PathoHR-Bench as a robust benchmark for evaluating a VL model’s ability to capture fine-grained pathological semantics. The case studies provided in the appendix D further illustrate the model’s performance on fine-grained diagnosis.
@@ 5.2 Comparison of existing VL models
Table 2: Zero-shot classification performance comparison using single prompt in terms of balanced accuracy and weighted F1 across four datasets. Best and second results are highlighted with bold and underline.
@@ 5.3 Ablation studies
Results are presented in the Table 3, which summarizes the individual contributions to model performance on PathoHR-Bench, along with average zero-shot accuracy across six pathology datasets. We observe that incorporating either negative text or negative images enhances structural awareness, but alone is insufficient to boost zero-shot performance. As detailed in Appendix E, the generation of negative samples introduces contrastive, parallel, and inclusion relationships; among them, inclusion relationships often lead to semantic confusion due to high similarity, which undermines contrastive learning effectiveness.
In contrast, positive text and images improve classification accuracy but may reduce the model’s sensitivity to structural variations. Our full scheme achieves optimal results by jointly enhancing structural understanding, compositional reasoning, and generalization ability. To further assess the clinical plausibility of the generated samples, we conducted a qualitative evaluation with a certified pathologist. The evaluation protocol and results are presented in Appendix F.
@@ 6 Conclusions
7 Limitations
Despite the improvements achieved with PathoHR-Bench and our proposed training scheme, several limitations remain. First, while the proposed benchmark systematically evaluates structural awareness and compositional reasoning, it does not yet incorporate external medical knowledge, which could further enhance model interpretability and diagnostic accuracy. Second, the perturbation strategies used in the framework, though effective, may not fully capture the diverse and nuanced variations present in real-world pathology reports.
@@ Appendix B Datasets
i). CRC100K (Kather et al., 2019): is a colorectal cancer dataset comprising $224\times 224$ pixel image patches acquired at $0.5\mu m$ per pixel resolution from 50 patients. It includes nine tissue categories: colorectal adenocarcinoma epithelium, normal mucosa, smooth muscle, lymphocytes, mucus, cancer-associated stroma, adipose tissue, background, and debris. The official split consists of 100,000 training and 7,180 testing images. For zero-shot tile-level classification, we directly evaluated on the testing split without any finetuning.
@@ Appendix C Implementation Details
We implemented all models using PyTorch and conducted training on 2 NVIDIA A100 GPUs. Across all vision-language pretraining variants, we used a temperature parameter of 0.02 for the contrastive loss and optimized using AdamW Loshchilov and Hutter (2017) with an initial learning rate of $5\times 10^{-6}$. A cosine decay scheduler was applied throughout training. All models were trained for 50 epochs with a batch size of 256. During fine-tuning, we adopted BioClinicalBERT (with a maximum token length of 512) as the text encoder, and PLIP-ViT-B/32-224 as the visual encoder. All images were resized to $224\times 224$, and standard data augmentations including random cropping, horizontal flipping, and color jittering were applied during training. For zero-shot classification, we used single prompts per class for evaluation, following the standard setup in prior works.
Appendix D Case Study
In this section, we present additional case studies to further demonstrate the improvement in fine-grained learning capabilities achieved by our proposed training scheme. Figure A1 illustrates examples of semantic shift perturbations at the connections level, information loss perturbations at the descriptors level, and order variation perturbations at the entities level within the PathoHR-Bench. Additionally, comparative results are presented on PLIP and our proposed training framework across these tasks, highlighting their performances in handling different types of perturbations. The results demonstrate that our proposed training scheme significantly enhances compositional reasoning and structural awareness, enabling the model to better understand pathological diagnostic logic and reasoning processes.
@@ Appendix D Case Study
Figure A1: Cases study of semantic shift perturbations at the connections level, information loss perturbations at the descriptors level, and order variation perturbations at the entities level within the PathoHR-Bench.
@@
Appendix E Pathological Relationships in Negative Sample Generation
@@
Appendix F Expert Evaluation of Generated Samples
@@ F.2 Evaluation Summary
F.3 Expert-Identified Limitations

### v2

@@ 1 Introduction
Recently, vision-language (VL) models have gained substantial advances (Goel et al., 2022; Li et al., 2022), fostering an integration of computer vision and natural language processing across a wide range of application domains (Wei et al., 2024; Das et al., 2024). In medical image analysis, contrastive learning-based approaches have been extensively employed to align large-scale medical images with corresponding diagnostic reports, thereby enabling zero-shot classification and grading without additional fine-tuning (Xie et al., 2024; Phan et al., 2024). This paradigm markedly reduces the reliance on high-quality annotated data and enhances the scalability of medical imaging models. Nevertheless, in contrast to other imaging modalities such as X-ray and MRI, pathological image analysis presents distinct and more formidable challenges due to its higher structural complexity and more subtle visual cues in the images.
Figure 1: Radar charts for compared models on PathoHR-Bench across multiple compositional reasoning aspects. The axes correspond to three types of perturbations: I (Information Loss), S (Semantic Drift), O (Order Variation), evaluated under three semantic roles: Entities, Descriptors and Connections. Higher values indicate stronger robustness.
Pathological biopsy microscopy remain the gold standard for tumor diagnosis, as it capture intricate cellular structures and morphological details (Ruiz López et al., 2017). However, the subtle intra-class variations in images pose significant challenges for disease classification. Existing approaches predominantly rely on task-specific models tailored to individual applications, such as Gleason grading of prostate cancer and breast cancer subtyping (Pati et al., 2023; Bulten et al., 2022). In contrast, pathological texts are inherently fine-grained, characterized by specialized medical terminology, precise pathological reasoning, and structured diagnostic logic. These narratives not only describe subtle visual cues but also offer critical insights for diagnostic grading and prognosis prediction.
@@ 2 Motivations
The lack of hierarchical reasoning capabilities significantly limits the reliability and clinical interpretability of existing VL models in pathological analysis. A further limitation lies in current evaluation practices, which predominantly rely on standard retrieval or classification metrics and offer limited insight into models’ structural understanding or reasoning ability. While recent studies in the natural image domain have proposed benchmarks for evaluating compositional reasoning in VL models (Zhao et al., 2022b; Diwan et al., 2022), these approaches are not directly applicable to pathology due to fundamental domain differences in textual structure and semantics. Descriptions in natural image datasets are typically open-domain and emphasize visually salient features. In contrast, pathological texts are “closed-domain,” relying on highly specialized and standardized vocabularies (e.g., MeSH, ICD-O) and structured diagnostic patterns. Unlike open-domain free text, these reports follow controlled terminology and precise reasoning logic, resulting in much tighter semantic relationships and long-range inferential chains.
@@ 3 Pathology Hierarchical Reasoning Benchmark (PathoHR-Bench)
At the text perturbation level, we have designed three core tasks to simulate the challenges encountered in real-world pathology texts, including information loss, semantic drift, and order variation:
i). Assessing sensitivity to information loss: We employ saliency-driven phrase deletion by leveraging the Pathology Language Image Pretraining (PLIP) (Huang et al., 2023) model to compute the similarity between each phrase and the corresponding image. The two most salient pathological terms are removed to create adversarial text variants, simulating the model’s reliance on critical textual cues and assessing its behavior under different deletion orders. It reflects how the model allocates information between visual and textual inputs while evaluating its robustness to local information losses.
ii). Assessing sensitivity to semantic drift: We employ LLM-guided token substitution, where a randomly selected word in each text is masked and then predicted using a pretrained medical BERT model (BioBERT) (Kenton and Toutanova, 2019; Lee et al., 2020). The generated adversarial samples test the model’s robustness to semantic drifts and fine-grained semantic comprehension. It evaluates whether the VL models can accurately differentiate between subtly different but diagnostically significant textual descriptions.
@@ 4.1 Pathology-guided textual perturbation
A simple but effective approach of negative text generation is to apply a collection of pre-defined linguistic rules to match and replace words associated with specific entity types or patterns (Doveh et al., 2023). We pre-defined seven pathological attribute dimensions (pathological states and grading, morphological features, histochemical characteristics, staining methods, anatomical structures and organs, biomolecular features, color information) via PubMed (Namata et al., 2012) and MeSH (Lipscomb, 2000), followed by random substitution at the level of pathological attributes to generate controlled perturbations.
@@ 4.5 Loss Function
Table 2: Zero-shot classification performance using a single prompt per class, reported as balanced accuracy and weighted F1 across six datasets. Best and second results are highlighted with bold and underline.
The total positive loss $L_{pos}$ is computed as:
$L_{pos}=L_{pos}^{text_{T}}+L_{pos}^{text_{I}}+L_{pos}^{img}$ (9)
Finally, the full fine-tuning loss of our proposed method can be written as:
$L=L_{con}+\alpha\cdot L_{neg}+\beta\cdot L_{pos}$ (10)
5 Experiments
5.1 Datasets
The ARCH dataset (Gamper and Rajpoot, 2021) is the only widely available pathology-specific paired image–text dataset, comprising 8,617 pairs extracted from pathology textbooks and PubMed research articles. To avoid overlap between training and evaluation, we carefully split ARCH into two disjoint subsets: textbook-derived samples were used to construct PathoHR-Bench, while PubMed-derived samples were reserved for training. Building on this resource, PathoHR-Bench was created by systematically applying three types of perturbations across three semantic roles. This expansion yielded a total of 77,553 image–text pairs for evaluation. The textbook portion provides more structured and hierarchical narratives, making it especially suitable for robust assessment of reasoning capabilities, while the PubMed portion ensures fair training for all baseline models.
We compared the performance of our proposed method with current pathological VL models, including baseline CLIP (Radford et al., 2021), PLIP (Huang et al., 2023), BiomedCLIP (Zhang et al., 2023), CONCH (Lu et al., 2024), QuiltNet (Ikezogwo et al., 2023), and MI-Zero (Lu et al., 2023c) on PathoHR-Bench. Inspired by the fair comparison strategy in CPLIP (Javed et al., 2024), we fine-tuned all baseline models on the ARCH dataset before evaluation, with the exception of CLIP. Notably, while PLIP was originally pre-trained on PubMed captions and various biomedical image–text pairs, it had not been fine-tuned on ARCH in its released version. To ensure consistent data distribution and minimize domain bias, we additionally fine-tuned PLIP and the other baselines on ARCH using their original loss objectives, without introducing our perturbation-based augmentations. CLIP remained a purely zero-shot baseline without fine-tuning on ARCH, serving as a general-purpose reference point. All models were then evaluated under the same testing splits and inference prompts for fair comparison.
For the zero-shot task, we utilized six publicly available datasets covering a range of cancer types, including four datasets at the patch level and two at the whole-slide level. To ensure fairness and consistency across all methods, we adopted a single prompt per class for evaluation, rather than using prompt ensembling. Detailed dataset descriptions are provided in the Appendix B, and implementation details can be found in the Appendix C. To disentangle the effect of ARCH-specific fine-tuning from the inherent representation ability of existing models under our unified single-prompt protocol, we additionally report the performance of BiomedCLIP and PLIP without any fine-tuning on ARCH in Appendix D.
5.2 Comparison of existing VL models
In contrast, our proposed method effectively leverages limited pathology-specific data to enable fine-grained compositional reasoning while maintaining strong performance on conventional classification tasks. As shown in Tables 1 and 2, the model achieves consistent improvements across nine textual perturbation settings and six zero-shot classification benchmarks. Notably, substantial gains are observed on CRC100K and UHU (with fine-grained cancer subtypes), while more moderate gains are seen on PanNuke and DigestPath (with coarser tumor/normal classification). These results further support PathoHR-Bench as a robust benchmark for evaluating a VL model’s ability to capture fine-grained pathological semantics. The case studies provided in the appendix E further illustrate the model’s performance on fine-grained diagnosis.
@@ 5.3 Ablation studies
Results are presented in the Table 3, which summarizes the individual contributions to model performance on PathoHR-Bench, along with average zero-shot accuracy across six pathology datasets. We observe that incorporating either negative text or negative images enhances structural awareness, but alone is insufficient to boost zero-shot performance. As detailed in Appendix F, the generation of negative samples introduces contrastive, parallel, and inclusion relationships; among them, inclusion relationships often lead to semantic confusion due to high similarity, which undermines contrastive learning effectiveness.
In contrast, positive text and images improve classification accuracy but may reduce the model’s sensitivity to structural variations. Our full scheme achieves optimal results by jointly enhancing structural understanding, compositional reasoning, and generalization ability. To further assess the clinical plausibility of the generated samples, we conducted a qualitative evaluation with a certified pathologist. The evaluation protocol and results are presented in Appendix G.
@@ 6 Conclusions
Potential Limitations
Despite the improvements achieved with PathoHR-Bench and our proposed training scheme, several limitations remain. First, while the proposed benchmark systematically evaluates structural awareness and compositional reasoning, it does not yet incorporate external medical knowledge, which could further enhance model interpretability and diagnostic accuracy. Second, although our negative sample construction was carefully designed to simulate realistic yet diagnostically challenging perturbations with both clinical plausibility and semantic precision, its scalability remains limited. Future work could explore more automated approaches, such as leveraging large language models to simplify perturbation generation and improve adaptability. Third, the perturbation strategies used in the framework, though effective, may not fully capture the diverse and nuanced variations present in real-world pathology reports.
@@ Appendix B Datasets
i). CRC100K (Kather et al., 2019): is a colorectal cancer dataset comprising $224\times 224$ pixel image patches acquired at $0.5\mu m$ per pixel resolution from 50 patients. It includes nine tissue categories: colorectal adenocarcinoma epithelium, normal mucosa, smooth muscle, lymphocytes, mucus, cancer-associated stroma, adipose tissue, background, and debris. The official split consists of 100,000 training and 7,180 testing images. For zero-shot tile-level classification, we directly evaluated on the testing split without any fine-tuning.
@@ Appendix C Implementation Details
We implemented all models using PyTorch and conducted training on 2 NVIDIA A100 GPUs. Across all vision-language pretraining variants, we used a temperature parameter of 0.02 for the contrastive loss and optimized using AdamW Loshchilov and Hutter (2017) with an initial learning rate of $5\times 10^{-6}$. A cosine decay scheduler was applied throughout training. All models were trained for 50 epochs with a batch size of 256. During fine-tuning, we adopted BioClinicalBERT (with a maximum token length of 512) as the text encoder, and PLIP-ViT-B/32-224 as the visual encoder. All images were resized to $224\times 224$, and standard data augmentations including random cropping, horizontal flipping, and color jittering were applied during training. For zero-shot classification, we used single prompts per class for evaluation, following the standard setup in prior works. Following a unified protocol, we finetune all pathology-focused baselines on the ARCH training split using their original objectives and without our perturbation-based augmentations, and then evaluate them on PathoHR-Bench and public datasets with a single prompt per class. Although PLIP is pretrained on large biomedical corpora, it is not originally finetuned on ARCH; we therefore additionally finetune PLIP on ARCH to reduce distribution shift. To preserve a general-purpose reference, CLIP remains a purely zero-shot baseline (no ARCH fine-tuning).
Appendix D Zero-shot without ARCH fine-tuning.
In Table A1, we observe that the performance differences between the settings with and without fine-tuning on ARCH are not very large, especially for PLIP. This is likely because PLIP was already pretrained on a sufficiently large number of pathology-specific image–text pairs, so plain fine-tuning on paired data may approach an upper limit for representation learning on coarser-grained tasks such as PanNuke and DigestPath. In contrast, our method yields clear improvements on fine-grained subtype classification (CRC100K and UHU) and on PathoHR-Bench even when trained with only limited ARCH data.
We attribute these gains to the use of carefully designed structured positives and targeted hard or soft negatives that explicitly supervise compositional reasoning over entities, descriptors, and relations, rather than relying solely on generic image-text alignment. Taken together, these results indicate that our framework can more effectively capture the inherent logic and hierarchical structure of pathology data beyond what standard fine-tuning can achieve.
Model (no ARCH finetune) UHU PanNuke DigestPath
Acc F1 Acc F1 Acc F1
BiomedCLIP (w/o finetune) 0.3371 0.1804 0.5019 0.5073 0.5481 0.5226
PLIP (w/o finetune) 0.3618 0.2085 0.6139 0.6014 0.7890 0.8027
Table A1: Zero-shot results without ARCH fine-tuning for two widely used baselines. Numbers are reported with a single prompt per class under our unified protocol.
Appendix E Case Study
In this section, we present additional case studies to further demonstrate the improvement in fine-grained learning capabilities achieved by our proposed training scheme. Figure A1 illustrates examples of semantic drift perturbations at the connections level, information loss perturbations at the descriptors level, and order variation perturbations at the entities level within the PathoHR-Bench. Additionally, comparative results are presented on PLIP and our proposed training framework across these tasks, highlighting their performances in handling different types of perturbations. The results demonstrate that our proposed training scheme significantly enhances compositional reasoning and structural awareness, enabling the model to better understand pathological diagnostic logic and reasoning processes.
@@ Appendix E Case Study
Figure A1: Cases study of semantic drift perturbations at the connections level, information loss perturbations at the descriptors level, and order variation perturbations at the entities level within the PathoHR-Bench.
@@
Appendix F Pathological Relationships in Negative Sample Generation
@@
Appendix G Expert Evaluation of Generated Samples
@@ G.2 Evaluation Summary
G.3 Expert-Identified Limitations